OpenStudy (anonymous):

If sin theta = 3/5 and theta has its terminal side in Quadrant II, find the exact value of tan 2theta.

OpenStudy (mertsj):



OpenStudy (mertsj):

\[\tan(2\theta)=\frac{2\tan \theta}{1-\tan^2\theta}=\frac{2(\frac{-3}{4})}{1-\frac{9}{16}}=\]

OpenStudy (anonymous):

So it would be equal to 24/7?

OpenStudy (mertsj):

\[\frac{\frac{-3}{2}}{\frac{7}{16}}=\frac{-24}{7}\]

OpenStudy (anonymous):

@Mertsj